# Diffusion State-Guided Projected Gradient for Inverse Problems

### v1

@@ 1 Introduction
Figure 1: High-level interpretation of Diffusion State-Guided Projected Gradient (DiffStateGrad). DiffStateGrad projects the measurement gradient onto a subspace, defined to capture statistics of the diffusion state at time $t$. This helps the process stay closer to the data manifold during the diffusion process, resulting in better posterior sampling. Without such projection, the measurement gradient pushes the process off the data manifold. The dotted straight line conceptually visualizes the subspace to which the measurement gradient is projected.
@@
Recent literature goes beyond such model-based priors and leverages information from data. The latest works employ generative diffusion models (Song & Ermon, 2019; Kadkhodaie & Simoncelli, 2021), which implicitly learn the data prior $p({\bm{x}})$ by learning a process that transforms noise into samples from a complex data distribution. For inverse problems, this reverse generation process is guided by the likelihood $p({\bm{y}}|{\bm{x}})$, forming a denoiser posterior, to generate data-consistent samples. While diffusion models show state-of-the-art performance, they still face challenges in solving inverse problems.
@@
The main challenge arises since the denoiser posterior, specifically the likelihood component $p({\bm{y}}|{\bm{x}})$, is intractable since the diffusion is trained unconditionally (Song et al., 2021). Prior work addresses this challenge by proposing various approximations or projections to the gradient related to the measurement likelihood $p({\bm{y}}|{\bm{x}})$ to achieve likely solutions (Kawar et al., 2022); when these approximations are not valid, it results in inaccurate posterior sampling and the introduction of “artifacts” in the reconstructed data (Chung et al., 2023). Latent diffusion models (LDMs) (Rombach et al., 2022), due to the nonlinearity of the latent-to-pixel decoder, further exacerbate this challenge. Besides this approximation, the lack of robustness of diffusion models to the measurement gradient step size (Peng et al., 2024) and the measurement noise, and the lack of guarantees for worst-case performance limits their practical applications for inverse problems.
@@
Our contributions: We propose a Diffusion State-Guided Projected Gradient (DiffStateGrad) to address the challenge of staying on the data manifold in solving inverse problems. We focus on gradient-based measurement guidance approaches that use the measurement as guidance to move the intermediate diffusion state ${\bm{x}}_{t}$ toward high-probability regions of the posterior. DiffStateGrad projects the measurement guidance gradient onto a low-rank subspace, capturing the learned prior at time $t$ in the diffusion process (Figure 1). We visualize how the diffusion process is pushed off the manifold when the measurement step size is relatively large in a diffusion model and how the incorporation of DiffStateGrad alleviates this challenge (Figure 2). We perform singular value decomposition (SVD) on the intermediate diffusion state of an image and use the $r$ highest contributing singular vectors as a choice of our projection matrix; by projecting the measurement gradient onto our proposed subspace, we aim to remove the directions orthogonal to the local manifold structure.
@@
  • We show that the crucial factor is the choice of the subspace, not the low-rank nature of the subspace projection. We find that our DiffStateGrad projection enhances performance, in contrast to random subspace projections and low-rank approximations of the measurement gradient (Table 1).
@@
  • We demonstrate that DiffStateGrad significantly improves the performance of state-of-the-art (SOTA) methods, especially in challenging tasks such as phase retrieval and high dynamic range reconstruction. For example, DiffStateGrad improves the PSNR of ReSample (Song et al., 2023a) from $27.61(8.07)$ to $31.19(4.33)$ for phase retrieval, reporting mean (std). Our experiments cover a wide range of linear inverse problems of box inpainting, random inpainting, Gaussian deblur, motion deblur, and super-resolution (Tables 2, 3 and 5) and nonlinear inverse problems of phase retrieval, nonlinear deblur, and high dynamic range (HDR) (Table 4) for image restoration tasks.
@@ 2 Background & Related Works
$d{\bm{x}}=-\frac{\beta_{t}}{2}{\bm{x}}\,dt+\sqrt{\beta_{t}}\,d{\bm{w}},$ (2)
@@
$d{\bm{x}}=\left[-\frac{\beta_{t}}{2}{\bm{x}}-\beta_{t}\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})\right]dt+\sqrt{\beta_{t}}d\bar{{\bm{w}}},$ (3)
@@
$d{\bm{x}}=\left[-\frac{\beta_{t}}{2}{\bm{x}}-\beta_{t}(\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}+\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t}))\right]dt+\sqrt{\beta_{t}}d\bar{{\bm{w}}}.$ (4)
@@
Figure 3: DiffStateGrad improves the worst-case performance. PSNR histogram of ReSample and DiffStateGrad-ReSample for phase retrieval shows that DiffStateGrad-ReSample has a significantly lower failure rate than ReSample.
Solving inverse problems with latent diffusion models. For complex scenarios where direct application of pixel-based models is computationally expensive or ineffective, latent diffusion models (LDMs) offer a promising alternative (Rombach et al., 2022). Given data ${\bm{x}}\in\mathbb{R}^{n}$, the LDM framework utilizes an encoder $\mathcal{E}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$ and a decoder $\mathcal{D}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$, with $k\ll n$, to work in a compressed latent space. ${\bm{x}}_{T}$ is encoded into a latent representation ${\bm{z}}_{T}=\mathcal{E}({\bm{x}}_{T})$ and serves as the starting point for the reverse diffusion process. Then, ${\bm{z}}_{0}$ is decoded to ${\bm{x}}_{0}=\mathcal{D}({\bm{z}}_{0})$, the final reconstruction of the image. Using a latent diffusion model introduces an additional complexity to solving inverse problems. The challenge arises from the nonlinear nature and non-uniqueness mapping of the encoder/decoder, discussed by Rout et al. (2023) in their proposed Posterior Sampling with Latent Diffusion (PSLD); PSLD improves performance by enforcing fixed-point properties on the representations.
@@
Prior works aim to address the first challenge by going beyond first-order statistics (Rout et al., 2024) or incorporating posterior covariance into the maximum likelihood estimation step (Peng et al., 2024). Other lines of work address the second issue by decoupling the measurement guidance from the sampling process; they update the data estimate $\hat{{\bm{x}}}_{0}$ at time $t$ using the measurement gradient before resampling it to the noisy manifold (Song et al., 2023a; Zhang et al., 2024). The above-discussed approaches are still highly sensitive to the measurement gradient step size (Peng et al., 2024). Indeed, balancing the measurement gradient with the unconditional score function remains a significant challenge to solving inverse problems using measurement-guided generation. Wu et al. (2024) avoids the discussed approximations and samples from the posterior directly to resolve the need to find a balance between measurement guidance and the prior process.
@@
Projections in diffusion models. Manifold and subspace projections are used in various contexts in diffusion models. He et al. (2024) uses a manifold-preserving approach to improve the efficiency of diffusion generation. Chung et al. (2022) proposes a manifold constraint to project the measurement gradient into the data manifold ${\bm{x}}_{0}$; our proposal is more effective, which projects the measurement gradient on the noisy diffusion state related to ${\bm{x}}_{t}$ or ${\bm{z}}_{t}$ to stay close to $\mathcal{M}_{t}$ rather than $\mathcal{M}_{0}$.
@@
Conditional diffusion models for inverse problems. We focus on unconditional diffusion models as learned priors to solve general inverse problems. This approach leverages already trained diffusion models, which is useful for domains with abundant data. Another approach is to train conditional diffusion models where $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}\mid{\bm{x}}_{t})$ is directly captured by the score function, or where the diffusion directly transforms the measurement into the underlying data (e.g., image-to-image diffusion) (Saharia et al., 2022; Liu et al., 2023; Chung et al., 2024). This latter approach is problem-specific; hence, it is not generalizable across inverse tasks.
@@ 3 Diffusion State-Guided Projected Gradient (DiffStateGrad)
We propose a Diffusion State-Guided Projected Gradient (DiffStateGrad) to solve general inverse problems. DiffStateGrad can be incorporated into a wide range of diffusion models to improve guidance-based diffusion models. Without loss of generality, we explain DiffStateGrad in the context of Latent-DPS (Chung et al., 2023) and note that DiffStateGrad applies to a wide range of pixel and latent diffusion-based inverse solvers (see Section 4).
@@
where $\bm{g}_{t}=\nabla_{{\bm{z}}_{t+1}}\log p({\bm{y}}\mid\hat{\bm{z}}_{0}({\bm{z}}_{t}))$ is the measurement gradient (MG), $\eta_{t}$ is the step size, and $\mathcal{P}_{\mathcal{S}_{t}}$ is a projection step onto the low-rank subspace $\mathcal{S}_{t}$. The main contribution of this paper is to define this subspace so it results in better posterior sampling; in other words, to define a subspace such that when the measurement gradient is projected onto, the diffusion process is not disturbed and pushed away from the data manifold. We show in Table 1 that indeed the subspace $\mathcal{S}_{t}$, defined by the intermediate diffusion state, results in an improved posterior sampling, unlike a subspace that is constructed based on a random matrix or the low-rank structure of the measurement gradient. Hence, we choose the diffusion state ${\bm{z}}_{t}$ to define $\mathcal{S}_{t}$.
@@
Let $\mathcal{M}$ be a smooth $m$-dimensional submanifold of a $d$-dimensional Euclidean space $\mathbb{R}^{d}$, where $m<d$. Assume that for each point ${\bm{z}}_{t}\in\mathcal{M}$, the tangent space $T_{{\bm{z}}_{t}}\mathcal{M}$ is well-defined, and the projection operator $\mathcal{P}_{\mathcal{S}_{{\bm{z}}_{t}}}$ onto an approximate subspace $\mathcal{S}_{{\bm{z}}_{t}}$ closely approximates the projection onto $T_{{\bm{z}}_{t}}\mathcal{M}$. For the state ${\bm{z}}_{t}\in\mathcal{M}$ and the measurement guidance gradient ${\bm{g}}_{t}\in\mathbb{R}^{d}$, consider two update rules:
@@
where $\eta>0$ is a small step size.
Then, for sufficiently small $\eta$, the projected update ${\bm{z}}_{t-1}^{\prime}$ stays closer to the manifold $\mathcal{M}$ than the standard update ${\bm{z}}_{t-1}$. That is,
@@
The remainder of this section provides intuitions on how DiffStateGrad improves solving inverse problems in the presence of a suitable learned prior. Let the initial latent state ${\bm{z}}_{t}$ be on the manifold $\mathcal{M}_{t}$ (e.g., being artifact-free, as the diffusion model is trained on clean data samples). We observe that pushing away from the manifold process (e.g., introducing artifacts) can only be introduced via the guidance by the data-consistency gradient step, as this is the sole mechanism by which information from the measurement process enters the latent space. Consider the manifold $\mathcal{M}$ of artifact-free latent representations. Each ${\bm{z}}_{t}$ lies on this manifold, and the tangent space $T_{{\bm{z}}_{t}}\mathcal{M}$ represents the directions of “allowable” updates that maintain the artifact-free property staying on the current manifold.
@@
Figure 2 demonstrates the effectiveness of DiffStateGrad in removing artifacts when the MG step size is large; artifacts are introduced onto the measurement gradient and stay within the latent representation in PSLD (Rout et al., 2023). On the other hand, the reverse process via DiffStateGrad-PSLD (our method applied to PSLD) stays artifact-free, consistent with the mathematical analysis in Proposition 1 and the practical efficacy of the proposed SVD-based subspace projection. Finally, we note that the most significant improvements appear in challenging tasks such as phase retrieval, HDR, and inpainting. We attribute the effectiveness of DiffStateGrad, particularly in challenging tasks, to a reduced rate of failure cases (Figure 3). By constraining solutions closer to the data manifold, DiffStateGrad minimizes extreme failures, enhances consistency in reconstruction quality, and improves overall performance metrics.
@@
3.1 Efficiency
DiffStateGrad introduces minimal computational overhead. We perform SVD at most once per iteration, and for latent diffusion solvers, this occurs in the latent space on $64\times 64$ matrices. By selecting a low rank based on a variance threshold, subsequent projection and reconstruction operations are performed on reduced matrices, further decreasing computational complexity. Figure 4 illustrates the runtime for the three diffusion-based methods of PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), and DAPS (Zhang et al., 2024) with and without DiffStateGrad. The figure shows that the additional computational cost of incorporating DiffStateGrad is marginal, typically adding only a few seconds to the total runtime (see Section C.2 for further details). We note that our method is not intended to improve efficiency, but rather to enhance performance and robustness.

### v2

@@ 1 Introduction
Figure 1: High-level interpretation of Diffusion State-Guided Projected Gradient (DiffStateGrad). DiffStateGrad projects the measurement gradient onto a subspace defined to capture statistics of the diffusion state at time $t$ on which the gradient guidance is applied. This helps the process stay closer to the data manifold during the diffusion process, resulting in better posterior sampling. Without such projection, the measurement gradient pushes the process off the data manifold. For when the measurement gradient guidance is applied to ${\bm{z}}_{0\mid t}$, the projection is defined to capture the structure of the tangent space of the clean data manifold. The dotted straight line conceptually visualizes the subspace to which the measurement gradient is projected.
@@
Recent literature goes beyond such model-based priors and leverages information from data. The latest works employ generative diffusion models (Song & Ermon, 2019; Kadkhodaie & Simoncelli, 2021), which implicitly learn the data prior $p({\bm{x}})$ by learning a process that transforms noise into samples from a complex data distribution. For inverse problems, this reverse generation process is guided by the likelihood $p({\bm{y}}|{\bm{x}})$, forming a denoiser posterior, to generate data-consistent samples. While diffusion models are state-of-the-arts, they still face challenges in solving inverse problems.
@@
The main challenge arises from the fact that the denoiser posterior, specifically the likelihood component $p({\bm{y}}|{\bm{x}})$, is intractable since the diffusion is trained unconditionally (Song et al., 2021). Prior work addresses this challenge by proposing various approximations or projections to the gradient related to the measurement likelihood $p({\bm{y}}|{\bm{x}})$ to achieve likely solutions (Kawar et al., 2022); when these approximations are not valid, it results in inaccurate posterior sampling and the introduction of “artifacts” in the reconstructed data (Chung et al., 2023). Latent diffusion models (LDMs) (Rombach et al., 2022), due to the nonlinearity of the latent-to-pixel deco der, further exacerbate this challenge. Besides this approximation, the lack of robustness of diffusion models to the measurement gradient step size (Peng et al., 2024) and the measurement noise, and the lack of guarantees for worst-case performance limits their practical applications for inverse problems.
@@
Our contributions: We propose a Diffusion State-Guided Projected Gradient (DiffStateGrad) to address the challenge of staying on the data manifold in solving inverse problems. We focus on gradient-based measurement guidance approaches that use the measurement as guidance to move the intermediate diffusion state ${\bm{x}}_{t}$ toward high-probability regions of the posterior. DiffStateGrad projects the measurement guidance gradient onto a low-rank subspace, capturing the data statistics of the learned prior (Figure 1). We visualize how the diffusion process is pushed off the manifold when the measurement step size is relatively large in a diffusion model and how the incorporation of DiffStateGrad alleviates this challenge (Figure 2). We define a projection step to preserve the measurement gradient on the tangent space of the state manifold. We achieve this projection by performing singular value decomposition (SVD) on the diffusion state of an image to which guidance is applied and use the $r$ highest contributing singular vectors as a choice of our projection matrix; by projecting the measurement gradient onto our proposed subspace, we aim to remove the directions orthogonal to the local manifold structure.
@@
  • We show that the crucial factor is the choice of the subspace, not the low-rank nature of the subspace projection. We find that our DiffStateGrad enhances performance. Our projection defines the subspace based on the structure of the state to which the measurement guidance is applied. This is in contrast to random subspace projections or low-rank approximations (Table 1).
@@
  • We demonstrate that DiffStateGrad significantly improves the performance of state-of-the-art (SOTA) methods, especially in challenging tasks such as phase retrieval and high dynamic range reconstruction. For example, DiffStateGrad improves the PSNR of ReSample (Song et al., 2023a) from $27.61(8.07)$ to $31.19(4.33)$ for phase retrieval, reporting mean (std). Our experiments cover a wide range of linear inverse problems of box inpainting, random inpainting, Gaussian deblur, motion deblur, and super-resolution (Tables 3(a) and 4(b)) and nonlinear inverse problems of phase retrieval, nonlinear deblur, and high dynamic range (HDR) (Table 3(b)) for image restoration tasks.
@@ 2 Background & Related Works
$d{\bm{x}}=-(\nicefrac{{\beta_{t}}}{{2}}){\bm{x}}\,dt+\sqrt{\beta_{t}}\,d{\bm{w}},$ (2)
@@
$d{\bm{x}}=[-(\nicefrac{{\beta_{t}}}{{2}}){\bm{x}}-\beta_{t}\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})]dt+\sqrt{\beta_{t}}d\bar{{\bm{w}}},$ (3)
@@
$d{\bm{x}}=[-(\nicefrac{{\beta_{t}}}{{2}}){\bm{x}}-\beta_{t}(\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}+\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t}))]dt+\sqrt{\beta_{t}}d\bar{{\bm{w}}}.$ (4)
@@
Figure 3: DiffStateGrad improves the worst-case performance. The PSNR histogram for phase retrieval shows that DiffStateGrad significantly lowers the failure rate.
Solving inverse problems with latent diffusion models. For complex scenarios where direct application of pixel-based models is computationally expensive or ineffective, latent diffusion models (LDMs) offer a promising alternative (Rombach et al., 2022). Given data ${\bm{x}}\in\mathbb{R}^{n}$, the LDM framework utilizes an encoder $\mathcal{E}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$ and a decoder $\mathcal{D}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$, with $k\ll n$, to work in a compressed latent space. ${\bm{x}}_{T}$ is encoded into a latent representation ${\bm{z}}_{T}=\mathcal{E}({\bm{x}}_{T})$ and serves as the starting point for the reverse diffusion process. Then, ${\bm{z}}_{0}$ is decoded to ${\bm{x}}_{0}=\mathcal{D}({\bm{z}}_{0})$, the final image reconstruction. Using a latent diffusion model introduces an additional complexity to solving inverse problems. The challenge arises from the nonlinear nature and non-uniqueness mapping of the encoder/decoder Rout et al. (2023); PSLD proposed to improve performance by enforcing fixed-point properties on representations.
@@
Prior works aim to address the first challenge by going beyond first-order statistics (Rout et al., 2024) or incorporating posterior covariance into the maximum likelihood estimation step (Peng et al., 2024). Other lines of work address the second issue by decoupling the measurement guidance from the sampling process; they update the data estimate $\hat{{\bm{x}}}_{0}$ at time $t$ using the measurement gradient guidance before resampling it to the noisy manifold at time $t-1$ (Song et al., 2023a; Zhang et al., 2024). The above-discussed approaches are still highly sensitive to the measurement gradient step size (Peng et al., 2024). Indeed, balancing the measurement gradient with the unconditional score function remains a significant challenge to solving inverse problems using measurement-guided generation. Wu et al. (2024) avoids the discussed approximations and samples from the posterior directly to resolve the need to find a balance between measurement guidance and the prior process.
@@
Projections in diffusion models. Manifold and subspace projections are used in various contexts in diffusion models. MPGD (He et al., 2024) uses a manifold-preserving approach to improve the efficiency of diffusion generation and solving inverse problems. While this method follows a similar sentiment as our proposed framework, it is only applicable when the measurement gradient is applied to $\hat{\bm{x}}_{0\mid t}$. Moreover, it requires the existence of an autoencoder for achieving manifold projection, and its performance is heavily dependent on the expressive power of the autoencoder. Unlike MPGD, DiffStateGrad is applicable to methods that apply the guidance to $\hat{\bm{x}}_{0\mid t}$/$\hat{\bm{z}}_{0\mid t}$ (i.e., ReSample and DAPS) and methods that apply the guidance to $\hat{\bm{x}}_{t}$/$\hat{\bm{z}}_{t}$ (i.e., PSLD and DPS). Chung et al. (2022a) proposes a manifold constraint to project the measurement gradient into the data manifold ${\bm{x}}_{0}$ while the guidance updates ${\bm{x}}_{t}$; our proposal is more effective since we project the measurement gradient on the noisy diffusion state related to ${\bm{x}}_{t}$ or ${\bm{z}}_{t}$, preserving the $t$ state on $\mathcal{M}_{t}$ rather than $\mathcal{M}_{0}$.
Method Gradient computation Gradient incorporation Projection
${\bm{x}}_{0\mid t}$/${\bm{z}}_{0\mid t}$ ${\bm{x}}_{t}$/${\bm{z}}_{t}$ ${\bm{x}}_{0\mid t}$/${\bm{z}}_{0\mid t}$ ${\bm{x}}_{t}$/${\bm{z}}_{t}$ ${\bm{x}}_{0\mid t}$/${\bm{z}}_{0\mid t}$
DPS
PSLD
ReSample
DAPS
MCG
MPGD
DiffStateGrad (ours)
Table 2: Gradient guidance computation, incorporation, and projection for diffusion-based inverse problems.
Gradient guidance incorporation.  Prior works differ from one another in two key aspects: (a) how the gradient is computed and (b) how the gradient is used to update the diffusion state. Table 2 categorizes prior works based on these characteristics. While a few approaches, such as diffusion-based MRI (Chung & Ye, 2022), compute the measurement gradient using ${\bm{x}}_{t}$, most recent literature has shifted toward using ${\bm{x}}_{0|t}$ for gradient computation. Regarding gradient incorporation, the literature is further subdivided. For instance, methods like DPS and PSLD use the measurement gradient to update the state at time $t$, whereas other approaches, such as ReSample, DAPS, and MPGD, apply the guidance to the conditional state at $0|t$ before resampling. Additionally, while the projections in MCG (Chung et al., 2022a) and MPGD (He et al., 2024) are restricted to ${\bm{x}}_{0|t}$, DiffStateGrad stands out by being applicable to both types of methods.
@@
Conditional diffusion models for inverse problems. We focus on unconditional diffusion models as learned priors to solve general inverse problems. This approach leverages already trained diffusion models, which is useful for domains with abundant data. Another approach is to train conditional diffusion models where $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}\mid{\bm{x}}_{t})$ is directly captured by the score function, or where the diffusion directly transforms the measurement into the underlying data (e.g., image-to-image diffusion) (Saharia et al., 2022; Liu et al., 2023; Chung et al., 2024). This latter approach is problem-specific; hence, it is not generalizable across inverse tasks. Finally, we note that while this work focuses on gradient-based guidance, prior work such as RePaint (Lugmayr et al., 2022) introduces a gradient-free masking strategy to solve inverse problems. Although RePaint is appealing, it is limited to inpainting tasks and scenarios where measurement noise is negligible.
@@ 3 Diffusion State-Guided Projected Gradient (DiffStateGrad)
We propose a Diffusion State-Guided Projected Gradient (DiffStateGrad) to solve inverse problems. DiffStateGrad can be incorporated into a wide range of diffusion models to improve guidance-based diffusion models. Without loss of generality, we explain DiffStateGrad in the context of Latent-DPS (Chung et al., 2023) which applies the measurement guidance to ${\bm{z}}_{t}$. We note that DiffStateGrad applies to a wide range of pixel and latent diffusion-based inverse solvers (see Section 4).
@@
where $\bm{g}_{t}=\nabla_{{\bm{z}}_{t+1}}\log p({\bm{y}}\mid\hat{\bm{z}}_{0}({\bm{z}}_{t}))$ is the measurement gradient (MG), $\eta_{t}$ is the step size, and $\mathcal{P}_{\mathcal{S}_{t}}$ is a projection step onto the low-rank subspace $\mathcal{S}_{t}$. The main contribution of this paper is a) to highlight that the measurement gradient should be projected onto a subspace imposed by the state being updated by the gradient (see gradient incorporation column in Table 2) and b) to define this subspace so it results in better posterior sampling; in other words, to define a subspace such that when the measurement gradient is projected onto, the diffusion process is not disturbed and pushed away from the data manifold. In Table 1, we show for PSLD that indeed the subspace $\mathcal{S}_{t}$, defined by the intermediate diffusion state, results in an improved posterior sampling, unlike a subspace that is constructed based on a random matrix or the low-rank structure of the measurement gradient. Hence, we choose the diffusion state ${\bm{z}}_{t}$ to define $\mathcal{S}_{t}$. Finally, for methods where the measurement gradient guidance is being applied to ${\bm{z}}_{0\mid t}$, we define the low-rank subspace based on ${\bm{z}}_{0\mid t}$ (see Table 2).
@@
Let $\mathcal{M}$ be a smooth $m$-dimensional submanifold of a $d$-dimensional Euclidean space $\mathbb{R}^{d}$, where $m<d$. Assume that for each state ${\bm{z}}_{t}\in\mathcal{M}$, the tangent space $T_{{\bm{z}}_{t}}\mathcal{M}$ is well-defined, and the projection operator $\mathcal{P}_{\mathcal{S}_{{\bm{z}}_{t}}}$ onto an approximate subspace $\mathcal{S}_{{\bm{z}}_{t}}$ closely approximates the projection onto $T_{{\bm{z}}_{t}}\mathcal{M}$. For the state ${\bm{z}}_{t}\in\mathcal{M}$ and measurement gradient ${\bm{g}}_{t}\in\mathbb{R}^{d}$, consider two update rules:
@@
where $\eta>0$ is a small step size. Then, for sufficiently small $\eta$, the projected update ${\bm{z}}_{t-1}^{\prime}$ stays closer to the manifold $\mathcal{M}$ than the standard update ${\bm{z}}_{t-1}$. That is,
@@
The remainder of this section provides intuition on how DiffStateGrad improves solving inverse problems in the presence of a suitable learned prior. Let the initial latent state ${\bm{z}}_{t}$ be on the manifold $\mathcal{M}_{t}$ (e.g., being artifact-free). The term “artifact-free” refers to the generation process of an unconditional diffusion model that is trained on clean data samples and provides an artifact-free trajectory from $\mathcal{M}_{T}$ to $\mathcal{M}_{0}$. We observe that pushing away from the manifold process (e.g., introducing artifacts) can only be introduced via the guidance by the data-consistency gradient step, as this is the sole mechanism by which information from the measurement process enters the latent space. Consider the manifold $\mathcal{M}$ of artifact-free latent representations. Each ${\bm{z}}_{t}$ lies on this manifold, and the tangent space $T_{{\bm{z}}_{t}}\mathcal{M}$ represents the directions of “allowable” updates that maintain the artifact-free property staying on the current manifold. Finally, we note that this motivates to project the gradient onto the tangent space of the data manifold where the guidance is applied. Alternatively, when the measurement gradient guidance is applied to ${\bm{z}}_{0|t}$, we define the projection step based on $\mathcal{S}_{0}$ (see also (He et al., 2024)).
@@
Figure 2 demonstrates the effectiveness of DiffStateGrad in removing artifacts when the MG step size is large; artifacts are introduced onto the measurement gradient and stay within the latent representation in PSLD (Rout et al., 2023). On the other hand, the reverse process via DiffStateGrad-PSLD (our method applied to PSLD) stays artifact-free, consistent with the mathematical analysis in Proposition 1 and the practical efficacy of the proposed SVD-based subspace projection. Finally, we note that the most significant improvements appear in challenging tasks such as phase retrieval, HDR, and inpainting. We attribute the effectiveness of DiffStateGrad, particularly in challenging tasks, to a reduced rate of failure cases (Figure 3). By constraining solutions closer to the data manifold, DiffStateGrad minimizes extreme failures, enhances consistency in reconstruction quality.
@@
Efficiency. DiffStateGrad introduces minimal computational overhead. We perform SVD at most once per iteration, and for latent diffusion solvers, this occurs in the latent space on $64\times 64$ matrices. By selecting a low rank based on a variance threshold, subsequent projection and reconstruction operations are performed on reduced matrices, further decreasing computational complexity. Figure 4 illustrates the runtime for the three diffusion-based methods of PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), and DAPS (Zhang et al., 2024) with and without DiffStateGrad. The figure shows that the additional computational cost of incorporating DiffStateGrad is marginal, typically adding only a few seconds to the total runtime (see C.2 for further details). We note that our method is not intended to improve efficiency, but rather to enhance performance and robustness.